OpenStudy (bringmethehorizon_):

Does this equal -1 or 1? 4m + 2(m + 1) = 9m + 5

OpenStudy (anonymous):

It equals -1

OpenStudy (bringmethehorizon_):

Thanks. :)

OpenStudy (anonymous):

No problem! :) we will never sleep

OpenStudy (bringmethehorizon_):

Hhahah. bmth

OpenStudy (anonymous):

4m+2(m+1)=9m+5 simplify 4m+2m+2=9m+5 combine like terms 6m+2=9m+5 solve for m -1=m